Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (anonymous):

suppose you flip a coin 100 times and get 87 heads and 13 tails. Based on these results, what is the probability that the next flip results in a head? If I have to answer in a decimal, isn't it .87?

OpenStudy (anonymous):

Answer: 1/2 Each flip has an independent probability of 1/2 I dont care if the first 1,000 flips are all heads. The chance the 1,001 st flip is a head is still 1/2.

OpenStudy (anonymous):

50:50

OpenStudy (anonymous):

0.50

OpenStudy (anonymous):

1/2 = 0.50

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!